The age of Betty's mother is 7 years more than twice Betty's age. The difference in their ages is 21 years. How old is Betty? Show the linear equation in one variable that you used then show each step in solving for your answer.
Answer by richwmiller(17219)

m=2b+7
m-b=21
m=b+21
2b+7=b+21
2b=b+14
b=14
m=35
check
35=2*14+7
35=28+7
35=35
ok